OpenStudy (anonymous):

Find the exact area of the circle with the given diameter. d = 12.6 cm A. 39.69 cm2 B. 158.76 cm2 C. 158.76π cm2 D. 39.69π cm2

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

\(A = \pi r^2\) Diameter is twice the radius, so 12.6 / 2 = 6.3 \(A = \pi 6.3^2\) Multiply 6.3 * 6.3 that is your answer in terms of pi.

OpenStudy (anonymous):

39.69

OpenStudy (anonymous):

how do i know if it has a π

OpenStudy (anonymous):

Yes, so our answer is \(36.69 \pi\)

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

does it have a π or cm2

OpenStudy (anonymous):

Because we never multiplied 'pi' \(A = \pi r^2\) The radius here is 6.3. \(A = \pi 6.3^2\) 6.3^2 is 39.69 \(A = 39.69 \pi\)

OpenStudy (anonymous):

Sorry I meant "Yes, so our answer is \(39.69 \pi\)"

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

Np